# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ 19             Break
We note, that although Algorithm 2 always provides a parent set of cardinality $m$, depending on the properties of the data set $\mathcal{X}_{n}$, the norming constant might be arbitrarily small. This is, however, inevitable. For some special, irregular data sets, its possible that none of the possible parent sets of $X_{i}$ are unisolvent and hence the norming constant $c_{N}$ is infinity. For a tight control on the norming constant, additional structural assumptions on the data set $\mathcal{X}_{n}$ is required. For example, the grid data (or its slightly perturbed version), investigated in the previous section provides such strong control.
B Proof of the lemmas and theorems in Section 5.1 and 5.2
B.1 Proof of Lemma 4
@@
By the definition of the Matérn covariance function in (2), for all multi-indices $k_{1},k_{2}$ satisfying $|k_{1}|+|k_{2}|\leq 2\underline{\alpha}$,
@@
Next we consider the derivative $K^{(k,k)}(x_{1},x_{2})$ for $|k|\leq\underline{\alpha}$, which can be directly computed as
@@
Noticing that the RHS of (47) (up to a constant multiplier) coincides with K(x,x)K(x,x+h)K(x,x)-K(x,x+h) for regularity parameter $\alpha-\underline{\alpha}$, it suffices to prove the upper bound (25) for $\alpha\in(0,1]$. From now on we will apply the representation (1) of the covariance kernel. For $\alpha\in(0,1)$, combining equations (10.25.2) and (10.27.4) of [41] yields the following expansion for the Matérn covariance kernel near zero
@@ B.2 Proof of Lemma 5
Next, using the notations $\langle A,\xi\rangle=(\langle w_{1},\xi\rangle,\langle w_{2},\xi\rangle,\ldots,\langle w_{m},\xi\rangle)^{T}$ and $A-x=(w_{1}-x,...,w_{m}-x)$, one can note that
@@
$\qquad\qquad+K_{x_{1},A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}(e^{-\iota\langle-A,\xi\rangle})^{T}K_{A,A}^{-1}K_{A,x_{2}}\Big{]}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi$
@@
$\qquad\qquad\times\overline{\big{(}e^{-\iota\langle x_{2},\xi\rangle}-K_{x_{2},A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}\big{)}}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi.$ (48)
@@
$g(x;\xi)=e^{-\iota\langle x,\xi\rangle}-K_{x,A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}.$ (49)
Noticing $K_{x,A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}$ is the Gaussian interpolation passing through $\{(x,e^{-\iota\langle x,\xi\rangle}),\forall x\in A\}$, the function $g(x;\xi)$ is $2\underline{\alpha}$ times differentiable and $g(x;\xi)=0,\forall x\in A$. Then, for all multi-index $k$, we denote its $k$th derivative as
@@
$\tilde{K}^{(k_{1},k_{2})}(x_{1},x_{2})=\int_{\mathbb{R}^{d}}g^{(k_{1})}(x_{1};\xi)\overline{g^{(k_{2})}(x_{2};\xi)}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi.$ (50)
When $x_{1}=x_{2}=x$ and $k_{1}=k_{2}=k$, equation (50) yields
$\tilde{K}^{(k,k)}(x,x)=\int_{\mathbb{R}^{d}}|g^{(k)}(x;\xi)|^{2}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi\geq 0.$
@@
In addition, following from (50),
$|\tilde{K}^{(k,k)}(x_{1},x_{2})|\leq$ $\int_{\mathbb{R}^{d}}\big{(}|g^{(k)}(x_{1};\xi)|^{2}+|g^{(k)}(x_{2};\xi)|^{2}\big{)}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi$
@@
Next we deal with (5). For all $|k|=\underline{\alpha}$ and $h\in\mathbb{R}^{d}$ in a neighborhood of zero, let us define an auxiliary process
@@
$\qquad\qquad-\tilde{K}^{(k,k)}(x_{1}+h,x_{2})+\tilde{K}^{(k,k)}(x_{1}+h,x_{2}+h).$ (51)
Thus it is sufficient to study the covariance function of the process $Y_{x}^{h}$. For $x_{1}=x_{2}=x$, we have
@@
$\mathrm{Cov}(Y_{x_{1}}^{h},Y_{x_{2}}^{h})\leq\max\{\mathrm{Var}(Y_{x_{1}}^{h}),\mathrm{Var}(Y_{x_{2}}^{h})\}\leq 2\sup_{x\in\mathbb{R}^{d}}\big{[}K^{(k,k)}(x,x)-K^{(k,k)}(x,x+h)\big{]}.$ (52)
The combination of (25), (B.2) and (52), provides our statement (5). ∎
B.3 Proof of Theorem 1
@@
The proof is organized in three steps. In the first and second steps, we fix the rescaling parameters $(\tau,s)$ as $(1,1)$ and provide matching lower and upper bounds. In the third step, we apply the rescaling parameters to finish the proof.
Direction $\gtrsim$: The case for $i=1$ is trivia. For all $i\geq 2$, we have
@@
Let us define the set $A_{i}:=\{X_{i}\}\cup\mathrm{pa}(X_{i})$. Then, applying Theorem 12.3 of [54] (with $\Phi$ being the Matérn covariance function, $\varphi_{0}(M)\asymp\inf_{\|\xi\|_{2}\leq 2M}(1+\|\xi\|_{2}^{2})^{-\alpha+d/2}=M^{-(2\alpha+d)}$, $q_{X}$ being $\gamma^{-j}$ and $M\asymp d/q_{X}$), the minimal eigenvalue of the matrix $K_{A_{i},A_{i}}$ satisfies that
@@
Direction $\lesssim$: Let us take $A=\mathrm{pa}(X_{i})$ in Lemma 5 and recall the formula for the derivative $\tilde{K}^{(k,k)}(x_{1},x_{2})$ in equation (50) and the definition of $g(x;\xi)$ in (49). Since $g$ is $2\underline{\alpha}$ times differentiable, recalling the definition of cube $\mathcal{C}$ in Condition 2, the $\underline{\alpha}$ order Taylor expansion of $g(x;\xi)$ around $x_{0}\in\mathcal{C}$ can be written as
$g(x;\xi)=Q(x;\xi)+R(x;\xi)$
where $Q(x;\xi)$ is a polynomial of order (up to) $\underline{\alpha}$ and $x\rightarrow R(x;\xi)$ is a $\underline{\alpha}$ times differentiable function with $R^{(k)}(x_{0};\xi)=0,\;\forall|k|\leq\underline{\alpha}.$ Thus for all multi-index $k$ satisfying $|k|=\underline{\alpha}$, the $k$th derivative of $Q(x;\xi)$ is a constant. Denoting this constant by $Q_{k}$ we get
$g^{(k)}(x;\xi)=Q_{k}+R^{(k)}(x;\xi).$ (53)
By plugging in equation (53) into equation (50) and the latter one further into the inequality (5), we get for all $\alpha\not\in\mathbb{N}$ and $|k|=\underline{\alpha}$,
@@
For all $x\in\mathcal{C}$, let $x=x_{0}+h$, then $\|h\|\lesssim\gamma^{-j}$ and in view of the previous display
$\int_{\mathbb{R}^{d}}|R^{(k)}(x;\xi)|^{2}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi\lesssim\gamma^{-2(\alpha-\underline{\alpha})j}.$ (54)
@@
$\int_{\mathbb{R}^{d}}|R^{(k)}(x;\xi)|^{2}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi\lesssim\gamma^{-2(\alpha-|k|)j}.$ (55)
@@
$\int_{\mathbb{R}^{d}}|R(x;\xi)|^{2}(1+\|\xi\|_{2}^{2})^{-(\alpha+d/2)}d\xi\lesssim\gamma^{-2\alpha j},\;\forall x\in\mathcal{C}.$ (56)
Recalling the definition of $g$ in (49) and the fact that $g(x;\xi)=0$ for all $x\in A\supset\mathrm{pa}(X_{i})$, we have $|Q(x;\xi)|=|R(x;\xi)|$. Furthermore, since $\mathrm{pa}(X_{i})$ is a norming set on the domain $\mathcal{C}$ for the collection of polynomials up to $\underline{\alpha}$,
@@
This in turn implies that for all $x\in\mathcal{C}$
@@
Then by plugging in the preceding upper bound into (50) and using (56), we arrive at
@@
Therefore, it only remained to prove assertion (55). We proceed via proof by induction on the decreasing order of $|k|$. The case $|k|=\underline{\alpha}$, is covered in (54). Now suppose (55) holds for all $k$ satisfying $|k|>l\in\mathbb{N}$. Denoting by $e_{j}\in\mathbb{N}^{d}$, $e_{j}[i]=\mathbbm{1}_{\{j=i\}}$ the elementary basis vector, we have for $|k|=l$,
@@
$\quad-R^{(k)}((x_{0}[1],\ldots,x_{0}[j],x[j+1],\ldots,x[d])^{T};\xi)\big{]}\bigg{|}^{2}(1+\|\xi\|_{2}^{2})^{-(\alpha+d/2)}d\xi$
@@
$\qquad\qquad\times(1+\|\xi\|_{2}^{2})^{-(\alpha+d/2)}d\xi$
@@
where in the first inequality we used Cauchy-Schwarz and in the second inequality Fubini’s theorem together with the induction assumption. This concoludes the proof of (55) and the upper bound for the statement for $s=\tau=1$.
Rescaling: For general scaling parameters $\tau$ and $s$ compared to the case $\tau=s=1$, the domain is mapped to $\tau\mathcal{X}$ and the covariance matrix is multiplied by a scalar $s^{2}$. Therefore by replacing $\gamma^{-j}$ with $\tau\gamma^{-j}$ and multiplying the final term with $s^{2}$, we obtain

### v4

@@ 19             Break
We note that, although Algorithm 2 always produces a parent set of cardinality $m$, the resulting norming constant may be arbitrarily small depending on the properties of the dataset $\mathcal{X}_{n}$. This limitation is, however, unavoidable. For certain irregular datasets, it is possible that none of the candidate parent sets for $X_{i}$ are unisolvent, in which case the norming constant $c_{N}$ becomes effectively infinity. To achieve a tight control over the norming constant, additional structural assumptions on the data set $\mathcal{X}_{n}$ are required. For example, grid data (or its slightly perturbed version), as considered in the previous section, provides such strong control.
D Computational aspects
In this section, we discuss the computational aspects of the posterior. We start by recalling the Bayesian hierarchical model introduced in Section 2.1. The observed data are $(X_{i},Y_{i})_{i=1,...,n}$, where $X_{i}$ are fixed design points in $0,1]^{d}$ and
The model parameters $f$ and $\sigma^{2}$ are endowed with a Vecchia Matérn GP and inverse-gamma prior, respectively, i.e.
$\sigma^{2}$ $\sim\mathrm{IG}(a_{0},b_{0}),$
$f$ $\sim(\hat{Z}_{X}^{\tau,s},X\in\mathcal{X}).$
Additionally, one can also introduce a hyper-prior on the scale parameters $\tau$ and $s$ (and/or the regularity parameter $\alpha$) of the Vecchia GP.
In the considered regression model with nuggets, i.e. observational errors $\epsilon_{i}$, the Cholesky factor of the posterior precision matrix is no longer sparse. As a result, we use a Gibbs sampling framework for inference. More concretely, we alternate sampling the variance parameter $\sigma^{2}$, the Vecchia GP $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}$, and any hyperparameters (if included) from their respective full conditional distributions.
Before providing further details, we introduce some additional notation. Let $\hat{\Phi}_{n}$ denote the precision matrix of the Vecchia GP $\hat{Z}_{X}^{\tau,s}$ evaluated on the training data $\mathcal{X}_{n}$, and let $\mathcal{Y}_{n}=[Y_{1},Y_{2},\cdots,Y_{n}]^{T}$. For a Vecchia GP defined in Section 2.2, its DAG structure (or equivalent, the parent sets $\mathrm{pa}(X_{i})$ for all $i$) uniquely determines a Cholesky decomposition of the precision matrix
$\hat{\Phi}=\hat{L}\hat{L}^{T}.$
Moreover, if the DAG is sparse, the Cholesky factor $\hat{L}$ is also sparse. For a detailed derivation, see Lemma S3 and its proof in [62].
Next, we discuss sampling $\sigma^{2}$, $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}$ and any hyperparameters from their full conditional distributions. By conjugacy the full conditional distribution of $\sigma^{2}$ is also inverse-gamma, allowing for direct sampling. The most challenging and computationally intensive part is sampling $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}|\mathcal{X}_{n},\mathcal{Y}_{n},\sigma^{2}$, which follows a multivariate $n$-dimensional Gaussian distribution
Note that the sparse Cholesky decomposition of $\hat{\Phi}$ does not directly induce a sparse Cholesky decomposition of $\hat{\Phi}+I_{n}/\sigma^{2}$. Instead, one can sample from this distribution by solving a linear inverse problem using conjugate gradient solver [41]. This approach only requires matrix-vector multiplications and can fully utilize the sparse structures of the Cholesky factor $\hat{L}$.
If the model includes hyper-priors on the GP parameters, e.g. scale or regularity, the corresponding hyper-posteriors are sampled using a Metropolis-Hasting step within Gibbs sampler. This requires evaluating both the prior density and the likelihood. The likelihood $\mathcal{Y}_{n}|\hat{Z}_{\mathcal{X}_{n}}^{\tau,s},\sigma^{2}$ is a product of independent Gaussian densities. The GP prior density for $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}$, in view of the sparse Choleksy decomposition of $\hat{\Phi}$, can be efficiently computed as
$p(\hat{Z}_{\mathcal{X}_{n}}^{\tau,s})=(2\pi)^{-n/2}\det(\hat{L})^{2}\exp\big[-(\hat{L}\hat{Z}_{\mathcal{X}_{n}}^{\tau,s})^{2}\big],$ (D.1)
where the determinant reduces to the product of the diagonal entries of $\hat{L}$ due to its triangular structure.
We summarize the full sampling scheme in Algorithm 3. Note that the method is fairly general and applies to any Vecchia GPs with a well-defined DAG structure – not only the layered norming DAGs and Matérn mother GPs proposed in this work.
1Input: training data $\mathcal{X}_{n},\mathcal{Y}_{n}$; regularity $\alpha$ and scale parameters $\tau,s$ of the GP (or their hyper-priors); number of MCMC iterations $n_{mc}$.
2 Build a DAG on the training locations $\mathcal{X}_{n}$.
3 for $1\leq i_{mc}\leq n_{mc}$ do
4   Sample $\sigma^{2}$ from its full conditional distribution;
5
6  Generate random vector ${W}\sim N(\mathcal{Y}_{n}/\sigma^{2},\hat{\Phi}+{I}_{n}/\sigma^{2})$ via ${W}=\mathcal{Y}_{n}/\sigma^{2}+\hat{{L}}{W}_{1}+{W}_{2}/\sigma,$ where ${W}_{1},{W}_{2}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,{I}_{n})$ and $\hat{L}$ is the Cholesky factor of the precision matrix $\hat{\Phi}$;
7   Obtain a sample of $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}$ by solving the linear system $(\hat{\Phi}+{I}_{n}/\sigma^{2})\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}={W}$ using the conjugate gradient method;
8   If hyper-priors are present, sample the hyperparameters, where the prior density of $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}$ is computed via (D.1).
9 end for
Result: Output the posterior samples of $\sigma^{2}$, $\hat{Z}_{\mathcal{X}_{n}}^{\tau,s}$, and hyperparameters (if applicable).
Algorithm 3 Posterior sampling for Vecchia GP
Computational complexity Considering a sparse DAG structure with parent set size bounded by $m$ – e.g., for the layered norming DAG introduced in Section 3, we have $m={\underline{\alpha}+d\choose\underline{\alpha}}$ – the computational complexity of Algorithm 3 is $O(nmn_{cg}n_{mc})$, where $n_{mc}$ denotes the number of MCMC iterations and $n_{cg}$ is the average number of conjugate gradient steps required per iteration. The term $nm$ corresponds to the cost of a single conjugate gradient step, as each matrix-vector multiplication involving the sparse Cholesky factor $\hat{L}$ can be carried out in $O(nm)$ time.
E Proof of the lemmas and theorems in Section 4.1 and 4.2
E.1 Proof of Lemma 1
@@
By the definition of the Matérn covariance kernel in (2), for all multi-indices $k_{1},k_{2}$ satisfying $|k_{1}|+|k_{2}|\leq 2\underline{\alpha}$,
@@
Next, we consider the derivative $K^{(k,k)}(x_{1},x_{2})$ for $|k|\leq\underline{\alpha}$, which can be directly computed as
@@
Noting that the right hand side (RHS) of (E.1) (up to a constant multiplier) coincides with K(x,x)K(x,x+h)K(x,x)-K(x,x+h) for regularity parameter $\alpha-\underline{\alpha}$, it is sufficient to prove the upper bound (20) for $\alpha\in(0,1]$. From now on we will apply the representation (1) of the covariance kernel. For $\alpha\in(0,1)$, combining equations (10.25.2) and (10.27.4) of [42] yields the following expansion for the Matérn covariance kernel near zero
@@ E.2 Proof of Lemma 2
Next, using the notation $\langle A,\xi\rangle=(\langle w_{1},\xi\rangle,\langle w_{2},\xi\rangle,\ldots,\langle w_{m},\xi\rangle)^{T}$ and $A-x=(w_{1}-x,...,w_{m}-x)$, one can note that
@@
$\qquad\qquad+K_{x_{1},A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}(e^{-\iota\langle-A,\xi\rangle})^{T}K_{A,A}^{-1}K_{A,x_{2}}\Big](1+\|\xi\|^{2}_{2})^{-(\alpha+d/2)}d\xi$
@@
$\qquad\qquad\times\overline{\big(e^{-\iota\langle x_{2},\xi\rangle}-K_{x_{2},A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}\big)}(1+\|\xi\|^{2}_{2})^{-(\alpha+d/2)}d\xi.$ (E.2)
@@
$g(x;\xi)=e^{-\iota\langle x,\xi\rangle}-K_{x,A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}.$ (E.3)
Noting that $K_{x,A}K_{A,A}^{-1}e^{-\iota\langle A,\xi\rangle}$ is the Gaussian interpolation passing through $\{(x,e^{-\iota\langle x,\xi\rangle}),\forall x\in A\}$, the function $g(x;\xi)$ is $2\underline{\alpha}$ times differentiable and $g(x;\xi)=0,\forall x\in A$. Then, for each multi-index $k$, we denote its $k$th derivative as
@@
$\tilde{K}^{(k_{1},k_{2})}(x_{1},x_{2})=\int_{\mathbb{R}^{d}}g^{(k_{1})}(x_{1};\xi)\overline{g^{(k_{2})}(x_{2};\xi)}(1+\|\xi\|^{2}_{2})^{-(\alpha+d/2)}d\xi.$ (E.4)
When $x_{1}=x_{2}=x$ and $k_{1}=k_{2}=k$, equation (E.4) yields
$\tilde{K}^{(k,k)}(x,x)=\int_{\mathbb{R}^{d}}|g^{(k)}(x;\xi)|^{2}(1+\|\xi\|^{2}_{2})^{-(\alpha+d/2)}d\xi\geq 0.$
@@
In addition, following from (E.4),
$|\tilde{K}^{(k,k)}(x_{1},x_{2})|\leq$ $\int_{\mathbb{R}^{d}}\big(|g^{(k)}(x_{1};\xi)|^{2}+|g^{(k)}(x_{2};\xi)|^{2}\big)(1+\|\xi\|^{2}_{2})^{-(\alpha+d/2)}d\xi$
@@
Next, we deal with (2). For all $|k|=\underline{\alpha}$ and $h\in\mathbb{R}^{d}$ in a neighborhood of zero, let us define an auxiliary process
@@
$\qquad\qquad-\tilde{K}^{(k,k)}(x_{1}+h,x_{2})+\tilde{K}^{(k,k)}(x_{1}+h,x_{2}+h).$ (E.5)
Thus, it is sufficient to study the covariance function of the process $Y_{x}^{h}$. For $x_{1}=x_{2}=x$, we have
@@
$\mathrm{Cov}(Y_{x_{1}}^{h},Y_{x_{2}}^{h})\leq\max\{\mathrm{Var}(Y_{x_{1}}^{h}),\mathrm{Var}(Y_{x_{2}}^{h})\}\leq 2\sup_{x\in\mathbb{R}^{d}}\big[K^{(k,k)}(x,x)-K^{(k,k)}(x,x+h)\big].$ (E.6)
The combination of (20), (E.2) and (E.6) provides our statement (2). ∎
E.3 Proof of Theorem 1
@@
The proof is organized in three steps. In the first and second steps, we fix the rescaling parameters $(\tau,s)=(1,1)$ and derive matching lower and upper bounds. In the third step, we apply the rescaling parameters to complete the proof.
Direction $\gtrsim$: The case $i=1$ is straightforward. For all $i\geq 2$, we obtain
@@
Define the set $A_{i}:=\{X_{i}\}\cup\mathrm{pa}(X_{i})$. Then, applying Theorem 12.3 of [59] (with $\Phi$ being the Matérn covariance function, $\varphi_{0}(M)\asymp\inf_{\|\xi\|_{2}\leq 2M}(1+\|\xi\|_{2}^{2})^{-\alpha+d/2}=M^{-(2\alpha+d)}$, $q_{X}$ being $\gamma^{-j}$ and $M\asymp d/q_{X}$), the minimal eigenvalue of the matrix $K_{A_{i},A_{i}}$ satisfies that
@@
Direction $\lesssim$: Take $A=\mathrm{pa}(X_{i})$ in Lemma 2 and recall the formula for the derivative $\tilde{K}^{(k,k)}(x_{1},x_{2})$ in equation (E.4) and the definition of $g(x;\xi)$ in (E.3). Since $g$ is $2\underline{\alpha}$ times differentiable, recalling the definition of the cube $\mathcal{C}$ in Condition 2, the $\underline{\alpha}$ order Taylor expansion of $g(x;\xi)$ around $x_{0}\in\mathcal{C}$ can be expressed as
$g(x;\xi)=Q(x;\xi)+R(x;\xi),$
where $Q(x;\xi)$ is a polynomial of order (up to) $\underline{\alpha}$ and $x\rightarrow R(x;\xi)$ is a $\underline{\alpha}$ times differentiable function with $R^{(k)}(x_{0};\xi)=0,\;\forall|k|\leq\underline{\alpha}.$ Thus, for any multi-index $k$ satisfying $|k|=\underline{\alpha}$, the $k$th derivative of $Q(x;\xi)$ is constant. Denoting this constant by $Q_{k}$, we get
$g^{(k)}(x;\xi)=Q_{k}+R^{(k)}(x;\xi).$ (E.7)
By plugging in equation (E.7) into equation (E.4) and the latter one further into the inequality (2), we get for all $\alpha\not\in\mathbb{N}$ and $|k|=\underline{\alpha}$,
@@
For all $x\in\mathcal{C}$, let $x=x_{0}+h$, then $\|h\|\lesssim\gamma^{-j}$ and in view of the previous display,
$\int_{\mathbb{R}^{d}}|R^{(k)}(x;\xi)|^{2}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi\lesssim\gamma^{-2(\alpha-\underline{\alpha})j}.$ (E.8)
@@
$\int_{\mathbb{R}^{d}}|R^{(k)}(x;\xi)|^{2}(1+\|\xi\|_{2})^{-(\alpha+d/2)}d\xi\lesssim\gamma^{-2(\alpha-|k|)j}.$ (E.9)
@@
$\int_{\mathbb{R}^{d}}|R(x;\xi)|^{2}(1+\|\xi\|_{2}^{2})^{-(\alpha+d/2)}d\xi\lesssim\gamma^{-2\alpha j},\;\forall x\in\mathcal{C}.$ (E.10)
Recalling the definition of $g$ in (E.3) and the fact that $g(x;\xi)=0$ for all $x\in A\supset\mathrm{pa}(X_{i})$, we have $|Q(x;\xi)|=|R(x;\xi)|$. Furthermore, since $\mathrm{pa}(X_{i})$ is a norming set on the domain $\mathcal{C}$ for the collection of polynomials up to degree $\underline{\alpha}$,
@@
This in turn implies that for all $x\in\mathcal{C}$,
@@
Then by plugging in the preceding upper bound into (E.4) and using (E.10), we arrive at
@@
Therefore, it only remains to prove assertion (E.9). We proceed by induction on the decreasing order of $|k|$. The case $|k|=\underline{\alpha}$ is covered in (E.8). Now suppose (E.9) holds for all $k$ satisfying $|k|>l\in\mathbb{N}$. Denoting by $e_{j}\in\mathbb{N}^{d}$, $e_{j}[i]=\mathbbm{1}_{\{j=i\}}$ the elementary basis vector, we have for $|k|=l$,
@@
$\qquad-R^{(k)}((x_{0}[1],\ldots,x_{0}[j],x[j+1],\ldots,x[d])^{T};\xi)\big]\bigg|^{2}(1+\|\xi\|_{2}^{2})^{-(\alpha+d/2)}d\xi$
@@
$\hskip 18.49988pt\hskip 18.49988pt\times(1+\|\xi\|_{2}^{2})^{-(\alpha+d/2)}d\xi$
@@
where in the first inequality we used Cauchy-Schwarz and in the second inequality we used Fubini’s theorem together with the induction assumption. This concludes the proof of (E.9) and the upper bound for the statement for $s=\tau=1$.
Rescaling: For general scaling parameters $\tau$ and $s$, compared to the case $\tau=s=1$, the domain is mapped to $\tau\mathcal{X}$ and the covariance matrix is scaled by a factor of $s^{2}$. Therefore, by replacing $\gamma^{-j}$ with $\tau\gamma^{-j}$ and multiplying the final term with $s^{2}$, we obtain